OpenStudy (anonymous):

Which of the sets of ordered pairs represents a function? A = {(3, −5), (4, 6), (−3, 9), (2, 7)} B = {(2, 4), (−1, −7), (5, 6), (4, 3)} Only A Only B Both A and B Neither A nor B

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

both

OpenStudy (anonymous):

How did you figure that out?

OpenStudy (anonymous):

there are no repeting inputs or x

OpenStudy (anonymous):

oh ok can you help me with another question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which equation represents a linear function? Equation 1: y = 2x + 1 Equation 2: y2 = 3x + 1 Equation 3: y = 5x5 − 1 Equation 4: y = 4x4 − 1 Equation 1 Equation 2 Equation 3 Equation 4

OpenStudy (anonymous):

Linear functions are those whose graph is a straight line. A linear function has one independent variable and one dependent variable. The independent variable is x and the dependent variable is y. a is the constant term or the y intercept.

OpenStudy (anonymous):

so A is the answer?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

is a, a line that goes strait on a graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):